OpenStudy (anonymous):

Simplify (3+2i)^2 5 - 12i 5 + 12i 9 - 4i 9 + 4i

HanAkoSolo (jamierox4ev3r):

Hint: (3+2i)^2=(3+2i)(3+2i)

HanAkoSolo (jamierox4ev3r):

based from this, can you solve? if not, let me know and i will teach you :D

OpenStudy (anonymous):

do you add or multiply?

OpenStudy (anonymous):

@Jamierox4ev3r

HanAkoSolo (jamierox4ev3r):

you would multiply

OpenStudy (anonymous):

9+4i..?

OpenStudy (anonymous):

i was factoring at first

OpenStudy (anonymous):

i^2 becomes -1

HanAkoSolo (jamierox4ev3r):

(3+2i)(3+2i) 9+6i+6i+4i^2

HanAkoSolo (jamierox4ev3r):

not necesarily @musiklover317

HanAkoSolo (jamierox4ev3r):

9+6i+6i+4i^2 9+12i+4i^2

OpenStudy (anonymous):

4i^2 = -4

HanAkoSolo (jamierox4ev3r):

here is the answer in simplest form: 4i^2+12i+9